# Accelerated Real-time Cine and Flow Under In-magnet Staged Exercise

### v2

@@ Abstract
Methods: This study investigated the feasibility of biventricular functional and hemodynamic assessment using real-time (RT) ExCMR during a staged exercise protocol in 24 healthy volunteers. We applied a coil reweighting technique and employed high acceleration rates to minimize motion blurring and artifacts. We further applied a beat-selection technique that identified beats from the end-expiratory phase to minimize the impact of respiration-induced through-plane motion. Additionally, results from six patients were presented to demonstrate clinical feasibility.
Results: Our findings indicated a consistent decrease in end-systolic volume and stable end-diastolic volume across exercise intensities, leading to increased stroke volume and ejection fraction. The selection of end-expiratory beats enhanced the repeatability of cardiac function parameters, as shown by scan-rescan tests in nine volunteers. High scores from a blinded image quality assessment indicated that coil reweighting effectively minimized motion artifacts.
@@
CMR, cardiovascular magnetic resonance imaging; CS, compressed sensing; LV, left ventricle; RV, right ventricle; EDV, end-diastolic volume; ESV, end-systolic volume; EF, ejection fraction; SV, stroke volume; CO, cardiac output; HR, heart rate; APHR, age-predicted maximal heart rate; NMAE, normalized mean absolute error; CCC, concordance correlation coefficient; AP, arbitrary respiratory phase; EE, end-expiratory phase; Vmax, peak velocity; RT, real-time; AAo, ascending aorta; MPA, main pulmonary artery; NFF, net forward flow; ExCMR, exercise stress CMR; ECG, electrocardiogram; BC, before correction; AC, after correction; BRPE, Borg rating of perceived exertion; UWT, undecimated wavelet transform; BBP, bean bag positioner; bSSFP, balanced steady-state free-precession; GRE, gradient echo; GRO, golden ratio offset; CAVA, Cartesian sampling with variable density and adjustable temporal resolution; R, acceleration rate; SCoRe, sparsity adaptive composite recovery; VENC, velocity encoding gradient
@@ 1 Introduction
Several studies have demonstrated the use of CMR immediately after in-room treadmill exercise [11]. In multiple clinical trials, this strategy has demonstrated success in detecting myocardial ischemia [12]. However, MRI-compatible treadmills are not readily available. Moreover, the need to transfer the patient from the treadmill to the MRI scanner not only creates a time delay between peak exercise and imaging but also makes imaging at various stress levels impractical. In contrast, the commercially available ergometers that enable supine exercise on the MRI patient table are more desirable. Several studies have shown the feasibility of out-of-bore exercise using ergometers [13, 14, 15], where the subject exercises on the MRI table outside the magnet. Once the target heart rate or exertion level is achieved, the subject is quickly moved into the scanner for imaging. Although more efficient than the treadmill protocol, this approach still suffers from some of the same limitations.
With advances in highly accelerated real-time (RT) imaging, ExCMR with in-magnet exercise has become an increasingly feasible and informative tool for assessing cardiac function under stress conditions. In 2009, Lurz et al. demonstrated the feasibility of accelerated RT ExCMR in 12 healthy subjects with an isotropic spatial resolution of 3 mm [16]. More recently, several other studies have shown the feasibility of RT ExCMR in small cohorts of healthy subjects [17, 18, 19] or patients [20, 15]. Collectively, these studies and the technical advances described within highlight the ongoing evolution of ExCMR. However, these studies still suffer from one or more of these limitations: (i) lower acceleration rates leading to poor in-plane spatial resolution ($>3$ mm), (ii) long scan times, (iii) smaller number of slices leading to limited coverage, (iv) dependence on electrocardiogram (ECG) which is typically not reliable during exercise, (v) lack of right ventricular quantification, (vi) absence of hemodynamic assessment, (vii) offline processing, (viii) lack of scan-rescan reproducibility, (ix) insufficient suppression of motion artifacts, especially under high-intensity exercise, (x) lack of a staged protocol, and (xi) sensitivity to through-plane motion.
In this study, we demonstrate the feasibility of biventricular hemodynamic and functional assessment of the heart using RT cine and flow data collected during a staged exercise protocol from 24 healthy subjects. More importantly, we present an in-magnet ExCMR protocol along with related data acquisition and processing techniques that address many of the technical limitations identified in recent ExCMR studies. To underline the clinical feasibility of RT ExCMR, we also present results from six patients.
@@ 2.1 Data acquisition
Twenty-eight healthy volunteers over the age of 18 were recruited to participate in an institutional review board-approved ExCMR study. Written informed consent was obtained from each subject. Data from three volunteers were excluded owing to their non-compliance with instructions or incidental findings. As summarized in Table 1, twenty-five volunteers were imaged on a 3T scanner (MAGNETOM Vida, Siemens Healthcare, Erlangen, Germany) fitted with an in-magnet supine ergometer (MR Ergometer Pedal, Lode, The Netherlands). The imaging protocol included a free-breathing RT short-axis cine stack (11 to 14 slices) covering the whole heart and at least two long-axis cine slices. Additionally, data were collected from two sets of three closely spaced phase-contrast CMR slices for measuring RT flow at the roots of the aortic and pulmonic arteries, as well as from 4D flow imaging with whole-heart coverage. The RT cine data were collected using a balanced steady-state free-precession (bSSFP) sequence, while the data for RT flow and 4D flow were collected using a gradient echo (GRE) sequence. For cine, a pseudo-random Cartesian sampling pattern, called golden ratio offset (GRO) [21], was used for prospective undersampling, while a different Cartesian sampling pattern, called Cartesian sampling with variable density and adjustable temporal resolution (CAVA) [22], was used for RT flow acquisition. GRO avoids large jumps in k-space between two consecutive readouts. This feature is important for avoiding eddy-currents-induced artifacts in bSSFP sequences [23]. In contrast, CAVA does not restrict the size of jumps in k-space but offers the added flexibility of changing the temporal resolution retrospectively. Since GRE sequences are less sensitive to large jumps in k-space, CAVA was used for RT flow. Both RT cine and flow acquisitions were highly accelerated, with the acceleration rates of $R\geq 8$ and $R\geq 16$, respectively. The protocol used to collect data is shown in Figure 1, and the imaging parameters for RT cine and flow are summarized in Table 2.
@@
After the ergometer was attached to the end of the table, each subject was positioned on the scanning table, with a medium-sized bean bag positioner (BBP) under their upper torso. The subjects were then asked to find a distance where they could pedal the ergometer comfortably without their knees touching the bore or over-extending their legs. Once a comfortable location was identified, their feet were secured to the ergometer pedals using Velcro straps and suction was applied to BBP using a vacuum device to create a firm mold around the upper torso. This mold helped the subjects stay in one position during in-magnetic exercise. Finally, a blood pressure cuff, a finger pulse oximeter, and ECG leads were attached to the subjects to monitor their vital signs during the scan.
After data acquisition at rest, the volunteers were asked to pedal the ergometer at a rate of 60 cycles per minute by matching their cadence to a metronome played through the headphones. The initial workload was set to $20$ W and was increased in increments of $20$ W to the maximum of $60$ W. The same set of sequences was executed at each exercise stage after allowing a ramp-up time of one minute for the heart rate (HR) to stabilize. Each exercise stage lasted for 8 to 10 minutes. Due to exhaustion or leg fatigue, not all subjects were able to complete all three exercise stages. In total, we had $n=25$, $n=24$, $n=23$, and $n=16$ volunteers at rest, $20$ W, $40$ W, and $60$ W, respectively. In one of the volunteers, an incidental finding during the resting stage of the protocol prompted termination of the experiment after the resting stage. The data from this volunteer was only used in the repeatability study. For a subset of volunteers, we repeated RT cine and flow acquisitions in quick succession (within 10 minutes) at rest ($n=11$) and at the $40$ W exercise stage ($n=9$) to assess repeatability. This study is focused on the analysis of RT cine and flow images. The data from the 4D flow acquisition will be reported separately [24].
@@
After completing the healthy subject study, six patient volunteers were also imaged on the same 3T scanner; see Table 1. In contrast to the healthy subjects, a $10$ W workload increment was employed when the $20$ W increase was considered excessively challenging by the patients. The maximum exercise intensity achieved by the six patients was $60$ W, $30$ W, $40$ W, $20$ W, $30$ W, and $20$ W.
@@ 2.2 Image reconstruction
The image reconstruction for RT cine and flow was performed inline with Gadgetron-based [26] implementation of a CS method, called sparsity adaptive composite recovery (SCoRe) [27]. A key feature of SCoRe is that it automatically adjusts the regularization strengths across multiple sparsifying transforms based on the image content. The data-driven and parameter-free nature of SCoRe makes it a great match for ExCMR, where temporal sparsity can change significantly with the level of exertion. Coil sensitivity maps were estimated using ESPIRiT [28]. In the case of RT cine, three-dimensional undecimated wavelet transform (UWT) was applied to sparsify both the spatial and temporal dimensions. For RT flow, in addition to UWT, temporal principal components [29], inferred after stacking flow-encoded and flow-compensated images, were used as data-driven sparsifying transform. The reconstruction was performed inline on a dedicated GPU workstation equipped with NVIDIA GeForce RTX 3090. For a $6$ s per slice acquisition, the reconstruction time was 10 seconds per slice for RT cine and 15 seconds per slice for RT flow.
Utilizing the BBP significantly limited the torso movement. In addition, the subjects were instructed to firmly hold onto the handles connected to the table’s rails during the exercise. Despite these measures, periodic bulk motion was observed during paddling in almost all cases. For some subjects, this resulted in significant degradation of image quality due to the temporally varying sensitivity maps of the body coil array. We attempted to use dynamic sensitivity maps, estimated from a sliding window of a small number of frames, but this approach was unsuccessful. Specifically, a smaller window (less than 8 frames) did not yield a fully sampled region of sufficient size or quality for reliable estimation of coil sensitivity maps, while a larger window (greater than 8 frames) introduced artifacts similar to those from time-averaged sensitivity maps. To mitigate these artifacts, we employed our recently proposed coil reweighting method [30], designed to automatically suppress contributions from artifact-inducing coil elements. This preprocessing step is computationally efficient and can be integrated with any reconstruction method, including CS.
@@ 2.3 Image analysis
For both RT cine and flow imaging, multiple consecutive heartbeats were collected from each slice. Several studies have highlighted the impact of respiration-induced variation in cardiac output quantification, which is attributed to changes in intrathoracic pressure and through-plane motion [31]. This variation is expected to be more pronounced during exercise. To minimize the impact of respiratory motion on cardiac function assessment, heartbeats from the end-expiratory phase were specifically isolated for analysis. The respiratory signal, critical for identifying this phase, was derived from RT images using a recently proposed principal component analysis-based method [32]. For imaging at rest, the ECG and respiratory signals facilitated the isolation of a single end-expiratory beat. Since ECG signal was unreliable during exercise, the extracted respiratory signal was superimposed and displayed at the bottom of the exercise RT image series. A moving white marker highlighted the respiratory location for each frame. This signal served as a reference to manually identify systolic and diastolic frames from the end-expiratory phase. An illustration is provided in Supporting Movie S1. The steps of isolating the end-expiratory beat (for rest images) and superimposing the respiratory signal (for exercise images) were performed offline.
@@
The DICOM images from the short-axis cine and flow acquisitions from the ascending aorta (AAo) and main pulmonary artery (MPA) were imported into suiteHEART (NeoSoft, Pewaukee, WI, USA) for hemodynamic and cardiac function quantification. For the RT data collected at rest, only one end-expiratory beat was imported. However, for RT data acquired during exercise stress, the entire cine and flow series, superimposed with the respiratory signal, were imported. Analyses were conducted only on expiratory heartbeats, aided by the reference respiratory signal. For an initial cohort of ten volunteers imaged with an acquisition time of 3 s/slice, all effort was made to conduct the analyses on the end-expiratory beat where possible. The protocol was consequently modified to an acquisition time of 6 s/slice to accommodate slower breathing patterns at rest. Biventricular segmentation was initially performed using suiteHEART software and manually adjusted where necessary. The following parameters were derived for cardiac function analysis: left ventricular (LV) and right ventricular (RV) end-diastolic volume (EDV), end-systolic volume (ESV), stroke volume (SV), and ejection fraction (EF). From the flow images, peak velocity (Vmax), net forward flow (NFF), and SV were measured for both major arteries. To account for the exaggerated through-plane motion, three RT flow slices were acquired from AAo and MPA. From the three slices available, a slice closest to but above the aortic valve was used for AAo flow quantification, and a slice closest to the pulmonic valve but away from the bifurcation was used for MPA flow analysis. The HR at rest and during each exercise stage was extracted from both RT cine and flow images and converted to a percentage of the age-predicted maximal heart rate (APHR). Cardiac output (CO), computed from both cine and flow images, was calculated as the product of SV (mL/beat) and HR (beats/min). In repeat imaging at rest and during a $40$ W workload, quantification was performed on a heartbeat from an arbitrary respiratory phase (AP) and on a heartbeat from end-expiration (EE).
To analyze the impact of coil reweighting on image quality, RT cine images from patients were scored with and without the reweighting procedure. From each subject, one mid-ventricular short-axis cine slice and two to three long-axis slices were scored by two expert readers on a five-point Likert-type scale (5-excellent, 4-good, 3-adequate, 2-fair, 1-non-diagnostic) in terms of overall image quality.

### v3

@@ Abstract
Methods: This study investigated the feasibility of biventricular functional and hemodynamic assessment using real-time (RT) ExCMR during a staged exercise protocol in 24 healthy volunteers. We employed high acceleration rates and applied a coil reweighting technique to minimize motion blurring and artifacts. We further applied a beat-selection technique that identified beats from the end-expiratory phase to minimize the impact of respiration-induced through-plane motion on cardiac function quantification. Additionally, results from six patients were presented to demonstrate clinical feasibility.
Results: Our findings indicated a consistent decrease in end-systolic volume and stable end-diastolic volume across exercise intensities, leading to increased stroke volume and ejection fraction. The selection of end-expiratory beats modestly enhanced the repeatability of cardiac function parameters, as shown by scan-rescan tests in nine volunteers. High scores from a blinded image quality assessment indicated that coil reweighting effectively minimized motion artifacts.
@@
AAo, ascending aorta; AP, arbitrary respiratory phase; APHR, age-predicted maximal heart rate; BBP, bean bag positioner; BMI, body mass index; BSA, body surface area; bSSFP, balanced steady-state free-precession; BRPE, Borg rating of perceived exertion; bpm, beats per minute; CAVA, Cartesian sampling with variable density and adjustable temporal resolution; CCC, concordance correlation coefficient; CMR, cardiovascular magnetic resonance imaging; CO, cardiac output; CR, coil reweighting; CS, compressed sensing; EDV, end-diastolic volume; ECG, electrocardiogram; EE, end-expiratory phase; EF, ejection fraction; ESV, end-systolic volume; ExCMR, exercise stress CMR; GRO, golden ratio offset; GRE, gradient echo; HR, heart rate; LV, left ventricle; MAD, mean absolute difference; MPA, main pulmonary artery; NFF, net forward flow; NMAE, normalized mean absolute error; R, acceleration rate; RT, real-time; RV, right ventricle; SCoRe, sparsity adaptive composite recovery; SV, stroke volume; UWT, undecimated wavelet transform; VENC, velocity encoding gradient; Vmax, peak velocity.
@@ 1 Introduction
Several studies have demonstrated the use of CMR immediately after in-room treadmill exercise [11]. In multiple clinical trials, this strategy has demonstrated success in detecting myocardial ischemia [12]. However, MRI-compatible treadmills are not readily available. Moreover, the need to transfer the patient from the treadmill to the MRI scanner not only creates a time delay between peak exercise and imaging but also makes imaging at various stress levels impractical. In contrast, the supine ergometers that enable exercise on the MRI patient table are more broadly available. Several studies have shown the feasibility of imaging during a pause in the in-magnet exercise [13, 14] or out-of-bore exercise using ergometers [15]. For the latter, the subject exercises on the MRI table outside the magnet. Once the target heart rate or exertion level is achieved, the subject is quickly moved into the scanner for imaging. Although more efficient than the treadmill protocol, these approaches still suffer from some of the same limitations.
With advances in highly accelerated real-time (RT) imaging, ExCMR during in-magnet exercise has become increasingly feasible. In 2009, Lurz et al. demonstrated the feasibility of accelerated RT ExCMR in 12 healthy subjects with an isotropic spatial resolution of 3 mm [16]. More recently, several other studies have shown the feasibility of RT ExCMR in small cohorts of healthy subjects [17, 18, 19] or patients [20, 21]. Collectively, these studies and the technical advances described within highlight the ongoing evolution of ExCMR. However, these studies still suffer from one or more of these limitations: (i) lower acceleration rates leading to poor spatial or temporal resolution, (ii) long scan times, (iii) smaller number of slices leading to limited coverage, (iv) dependence on electrocardiogram (ECG) which is typically not reliable during exercise, (v) lack of right ventricular quantification, (vi) absence of hemodynamic assessment, (vii) offline image reconstruction, (viii) lack of demonstration for scan-rescan reproducibility, (ix) insufficient suppression of motion artifacts, especially under high-intensity exercise, (x) lack of a staged protocol, and (xi) sensitivity to through-plane motion.
In this study, we demonstrate the feasibility of biventricular hemodynamic and functional assessment of the heart using RT cine and flow data collected during a staged exercise protocol from 24 healthy subjects. More importantly, we present an in-magnet ExCMR protocol and related data acquisition and processing techniques, including coil reweighting to suppress motion artifacts, that address many of the technical limitations identified in recent ExCMR studies. To underline the clinical feasibility of RT ExCMR, we also present results from six patients.
@@ 2.1 Data acquisition
Twenty-eight healthy volunteers over the age of 18 were recruited to participate in an institutional review board-approved ExCMR study. Written informed consent was obtained from each subject. Three volunteers were excluded owing to their non-compliance with instructions or incidental findings. As summarized in Table 1, twenty-five volunteers were imaged on a 3T scanner (MAGNETOM Vida, Siemens Healthcare, Erlangen, Germany) fitted with an in-magnet supine ergometer (MR Ergometer Pedal, Lode, The Netherlands). The imaging protocol included a free-breathing RT short-axis cine stack (11 to 14 slices) covering the whole heart and at least two long-axis cine slices. Additionally, phase-contrast CMR data were collected from two sets of three closely spaced slices for measuring RT flow at the roots of the aortic and pulmonic arteries, as well as from 4D flow imaging with whole-heart coverage. The RT cine data were collected using a balanced steady-state free-precession (bSSFP) sequence, while the data for RT flow and 4D flow were collected using a gradient echo (GRE) sequence. For cine, a pseudo-random Cartesian sampling pattern, called golden ratio offset (GRO) [22], was used for prospective undersampling, while a different Cartesian sampling pattern, called Cartesian sampling with variable density and adjustable temporal resolution (CAVA) [23], was used for RT flow acquisition. GRO avoids large jumps in k-space between two consecutive readouts. This feature is important for avoiding eddy-currents-induced artifacts in bSSFP sequences [24]. In contrast, CAVA does not restrict the size of jumps in k-space but offers the added flexibility of changing the temporal resolution retrospectively. Since GRE sequences are less sensitive to large jumps in k-space, CAVA was used for RT flow. To maintain adequate spatial and temporal resolutions, both RT cine and flow acquisitions were highly accelerated, with acceleration rates of $R\geq 8$ and $R\geq 16$, respectively. The exercise protocol used to collect data is shown in Figure 1, and the imaging parameters for RT cine and flow are summarized in Table 2.
@@
After the ergometer was attached to the end of the table, each subject was positioned on the scanning table, with a medium-sized bean bag positioner (BBP) under their upper torso. The subjects were instructed to find a comfortable position where they could pedal the ergometer without their knees touching the bore or over-extending their legs. Then, their feet were secured to the ergometer pedals using Velcro straps and suction was applied to BBP using a vacuum device to create a firm mold around the upper torso. This mold helped maintain the subject’s position during in-magnet exercise. Finally, a blood pressure cuff, a finger pulse oximeter, and ECG leads were attached to the subjects to monitor their vital signs during the scan.
After data acquisition at rest, the volunteers were asked to pedal the ergometer at a rate of 60 cycles per minute by matching their cadence to a metronome played through the headphones. The initial workload was set to $20$ W and was increased in increments of $20$ W to the maximum of $60$ W. The same set of sequences was executed at each exercise stage after allowing a ramp-up time of one minute for the heart rate (HR) to stabilize. Each exercise stage lasted for 8 to 10 minutes. Due to exhaustion or leg fatigue, not all subjects were able to complete all three exercise stages. In total, we scanned $n=25$, $n=24$, $n=23$, and $n=16$ volunteers at rest, $20$ W, $40$ W, and $60$ W, respectively. In one of the volunteers, an incidental finding prompted the termination of the experiment after the resting stage. The data from this volunteer was only used in the repeatability study. For a subset of volunteers, we repeated RT cine and flow acquisitions in quick succession (within 2 minutes) at rest ($n=11$) and at the $40$ W exercise stage ($n=9$) to assess repeatability. The order of acquisition was RT cine, RT flow, RT cine (repeat), RT flow (repeat), and 4D flow. This study is focused on the analysis of RT cine and flow images; the data from the 4D flow acquisition will be reported separately [25].
@@
After completing the healthy subject study, six patient volunteers were also imaged on the same 3T scanner; see Table 1. Four patients had a history of unexplained dyspnea on exertion, and two patients had asymptomatic moderate valvular heart disease (one moderate aortic stenosis and one moderate mitral regurgitation). In contrast to the healthy subjects, a $10$ W workload increment was employed whenever the $20$ W increase was considered excessively challenging by the patient. The maximum exercise intensity achieved by the six patients was $60$ W, $30$ W, $40$ W, $20$ W, $30$ W, and $20$ W.
2.2 Coil reweighting
Utilizing the BBP significantly limited the torso movement. In addition, the subjects were instructed to firmly hold onto the handles connected to the table’s rails during the exercise. Despite these measures, periodic bulk motion was observed during paddling in almost all cases. For some subjects, this resulted in significant degradation of image quality due to the physical movement of some of the receive coils resulting in temporally varying sensitivity maps. We attempted to use dynamic sensitivity maps, estimated from a sliding window of a small number of frames, but this approach was unsuccessful. Specifically, a smaller window (less than 8 frames) did not yield a fully sampled region of sufficient size or quality for reliable estimation of coil sensitivity maps, while a larger window introduced artifacts similar to those from time-averaged sensitivity maps.
To address this issue, we propose the pipeline shown in Figure 2A to automatically suppress the contribution from coils that introduce most of the artifacts. We first average the k-space of all the frames to generate the time-averaged k-space for each of the $N$ coils. The coil sensitivity maps $\{\boldsymbol{s}_{i}\}_{i=1}^{N}$ are then estimated using ESPIRiT [27], and the time-averaged coil images $\{\boldsymbol{x}_{i}\}_{i=1}^{N}$ are derived by performing inverse Fourier transformation of the time-averaged k-space. Subsequently, the time-averaged coil images are first combined and then re-estimated using the coil sensitivity maps, i.e.,
$\tilde{\boldsymbol{x}}_{i}=\left(\sum_{j=1}^{N}\boldsymbol{x}_{j}\odot~{}\boldsymbol{s}_{j}^{*}\right)\odot\boldsymbol{s}_{i},~{}\text{for~{}}i=1,2,\dots,N,$ (1)
where, $\tilde{\boldsymbol{x}}_{i}$ is the re-estimated time-averaged image from the $i^{\text{th}}$ coil, and $\odot$” and “*” represent element-wise multiplication and conjugation, respectively.
Considering that the ESPIRiT coil sensitivity maps are normalized, i.e., $\sum_{j=1}^{N}\boldsymbol{s}_{j}\odot\boldsymbol{s}_{j}^{*}$ returns an array of all-ones, one should expect $\tilde{\boldsymbol{x}}_{i}=\boldsymbol{x}_{i}$ in Equation (1) under the ideal condition of $\boldsymbol{x}_{i}=\overline{\boldsymbol{x}}\odot\boldsymbol{s}_{i}$, where $\overline{\boldsymbol{x}}$ represents the true coil-combined image. We propose using the residual signal, $\|\tilde{\boldsymbol{x}}_{i}-\boldsymbol{x}_{i}\|_{2}$, to identify coil elements that have inconsistent data due to their motion. After computing $\|\tilde{\boldsymbol{x}}_{i}-\boldsymbol{x}_{i}\|_{2}$ for each coil, we select $M$ coils with the largest $\|\tilde{\boldsymbol{x}}_{i}-\boldsymbol{x}_{i}\|_{2}$ values. The signal for each of those $M$ coils is scaled (reweighted) with $C^{2}/\|\tilde{\boldsymbol{x}}_{i}-\boldsymbol{x}_{i}\|_{2}^{2}$, where $C$ represents the baseline value computed by averaging residuals from $K$ coils with the smallest residual values. The weighting for the remaining $N-M$ coils is left unchanged. In this work, we chose $M=10$ and $K=15$. For a RT cine image series collected during exercise, Figure 2B shows the time-averaged and re-estimated images from two physical coils, one with significant motion artifacts (Coil-4) and the other without (Coil-11). The difference between $\boldsymbol{x}_{i}$ and $\tilde{\boldsymbol{x}}_{i}$ can be mainly attributed to motion artifacts. Figure 2C shows $\|\tilde{\boldsymbol{x}}_{i}-\boldsymbol{x}_{i}\|_{2}$ for all physical coils.
@@ 2.3 Image reconstruction
The coil reweighting was implemented inline as a pre-processing step and applied to all RT cine data. After coil reweighting, we compressed the k-space data to $12$ virtual coils for faster processing. The image reconstruction for RT cine and flow was also performed inline with Gadgetron-based [28] implementation of a CS method, called sparsity adaptive composite recovery (SCoRe) [29]. A key feature of SCoRe is that it automatically adjusts the regularization strengths across multiple sparsifying transforms based on the image content. The data-driven and parameter-free nature of SCoRe makes it well-suited for ExCMR, where temporal sparsity can change significantly with the level of exertion. The coil sensitivity maps of the 12 virtual coils were estimated using ESPIRiT [27]. In the case of RT cine, three-dimensional undecimated wavelet transform (UWT) was applied to sparsify both the spatial and temporal dimensions. For RT flow, in addition to UWT, temporal principal components [30], inferred after stacking flow-encoded and flow-compensated images, were used as data-driven sparsifying transform. The reconstruction was performed on a dedicated GPU workstation equipped with NVIDIA GeForce RTX 3090. For a $6$ s per slice acquisition, the reconstruction time was 10 seconds per slice for RT cine and 15 seconds per slice for RT flow.
@@ 2.4 Image analysis
For cardiac function quantification from RT cine and flow, multiple consecutive heartbeats were available from each slice. Several studies have highlighted the impact of respiration-induced variation in cardiac output quantification, which is attributed to changes in intrathoracic pressure [31]. Also, failing to account for the respiratory motion when selecting cardiac beats from different slices can introduce additional variability due to the through-plane motion. With exercise, this variability is expected to be more pronounced due to exaggerated breathing. To minimize the impact of respiratory motion on cardiac function assessment, heartbeats from the end-expiratory phase were specifically isolated for analysis. The respiratory signal, critical for identifying this phase, was derived from RT images using a recently proposed principal component analysis-based method [32]. For imaging at rest, the ECG and respiratory signals facilitated the isolation of a single end-expiratory beat. Since the ECG signal was unreliable during exercise, the extracted respiratory signal was superimposed and displayed at the bottom of the exercise RT image series. A moving white marker highlighted the respiratory location for each frame. This signal served as a reference to manually identify systolic and diastolic frames from the end-expiratory phase. An illustration is provided in Supporting Movie S1. The steps of isolating the end-expiratory beat (for rest images) and superimposing the respiratory signal (for exercise images) were performed offline.
@@
The DICOM images from the short-axis cine and flow acquisitions from the ascending aorta (AAo) and main pulmonary artery (MPA) were imported into suiteHEART (NeoSoft, Pewaukee, WI, USA) for hemodynamic and cardiac function quantification. For the RT data collected at rest, only one end-expiratory beat was imported. However, for RT data acquired during exercise stress, the entire cine and flow series, superimposed with the respiratory signal, were imported. Analyses were conducted only on expiratory heartbeats, aided by the reference respiratory signal. For an initial cohort of ten volunteers imaged with an acquisition time of 3 s/slice, every effort was made to conduct the analyses on the end-expiratory beat where possible. The protocol was consequently modified to an acquisition time of 6 s/slice to accommodate slower breathing patterns at rest. Biventricular segmentation was initially performed using suiteHEART software and manually adjusted where necessary. The following parameters were derived for cardiac function analysis: left ventricular (LV) and right ventricular (RV) end-diastolic volume (EDV), end-systolic volume (ESV), stroke volume (SV), and ejection fraction (EF). From the flow images, peak velocity (Vmax), net forward flow (NFF), and SV were measured for both major arteries. To account for the exaggerated through-plane motion, three RT flow slices were acquired from AAo and MPA. From the three slices available, a slice closest to but above the aortic valve was used for AAo flow quantification, and a slice closest to the pulmonic valve but away from the bifurcation was used for MPA flow analysis. The HR at rest and during each exercise stage was extracted from both RT cine and flow images and converted to a percentage of the age-predicted maximal heart rate (APHR). Cardiac output (CO), computed separately from cine and flow images, was calculated as the product of SV (mL/beat) and HR (beats/min). The values of EDV, ESV, SV, NFF, and CO were indexed using body surface area. In repeat imaging at rest and during a $40$ W workload, quantification was performed on a heartbeat from an arbitrary respiratory phase (AP) and on a heartbeat from end-expiration (EE). The repeatability was assessed using concordance correlation coefficient (CCC) and normalized mean absolute error (NMAE). NMAE, inferred from the two repeats, $x_{1}$ and $x_{2}$, was calculated as $\mathrm{NMAE}=\frac{100}{L}\times\sum_{i=1}^{L}\left[(|x_{1,i}-x_{2,i}|)/(\frac{1}{2}|x_{1,i}+x_{2,i}|)\right]$, where $L$ is the number of paired measurements.
To analyze the impact of coil reweighting on image artifacts, the RT-cine series reconstructed with and without coil reweighting were visually scored by two cardiologists for the level of artifacts. Ten healthy volunteers were randomly chosen along with all six patients. From each of these 16 subjects, one mid-ventricular short-axis slice and one two-chamber slice were selected at rest and the maximum level of exertion. The cine series, displayed as movies, were evaluated using a five-point scale: 1–Unusable, 2–Severe artifacts obscuring useful information, 3–Moderate artifacts with some loss of information but still diagnostic, 4–Minor artifacts with minimal loss of information, 5–No artifacts. Since the flow images exhibited minimal motion artifacts, coil reweighting was not applied to them. The motion artifacts in the bSSFP-based cine images are more pronounced because the inaccurate sensitivity maps cause the signal from the bright subcutaneous fat to alias to the rest of the image.
To statistically assess the impact of coil reweighting, paired t-tests were conducted using a significance level of 0.05. To account for potential non-independence, we adopted a conservative approach by averaging scores across two views and two readers prior to performing the t-tests. Although Likert-type data are ordinal, the t-test was employed under the assumption that the averaged scores approximate interval-level data. Additionally, inter-reader agreement for artifact scores was evaluated using the weighted Kappa statistic.